the endpoints of a line segment ab are a(3,-4) and b(13,d). the midpoint of ab is m (m,4). find the values of d and m and the equation of the line ab.
Answer by amarjeeth123(569)

the endpoints of a line segment ab are a(3,-4) and b(13,d)
The midpoint of ab=((13+3)/2,(d-4)/2)=(8,(d-4)/2)
the midpoint of ab is m (m,4)
m=8
(d-4)/2=4
d-4=8
d=12
Answer:m=8 and d=12